# Reevaluating Policy Gradient Methods for Imperfect-Information Games

### v1

@@ 4 Benchmarks
Unfortunately, evaluating DRL algorithms for adversarial IIGs is not well standardized. Due to the difficulty of efficiently implementing exploitability for the large IIGs, there is a dearth of accessible implementations. As a result, only a small number of works (dream20; armac20; nfsp-eve23) make the effort to report exploitability in large games.
The rest of literature predominantly resorts to some combination of two unsatisfactory metrics. The first is exploitability in the same small EFGs that served as benchmarks for tabular solvers, such as Leduc Hold’em (leduc05). Benchmarking in these small games is unsatisfactory because success requires learning a precise near-Nash policy for a handful of repeatedly revisited information states—a fundamentally different challenge from large games, where success requires learning a policy that is strong (but not necessarily numerically close to Nash) for a vast number of unvisited information states. The second is head-to-head evaluations in large games. These evaluations are unsatisfactory because the intransitive dynamics typically present in imperfect-information games, such as rock-beats-scissors-beats-paper-beats-rock, muddle interpretations of relative strength. Exacerbating the issue, there is no community coordination on head-to-head opponents, making direct cross-paper comparisons impossible.
@@
To address this predicament, we set out to implement efficient computations for four large games—Phantom Tic-Tac-Toe (PTTT), 3x3 Dark Hex (DH3), Abrupt Phantom Tic-Tac-Toe (APTTT), and 3x3 Abrupt Dark Hex (ADH3)—ourselves. These games are imperfect-information twists on the perfect-information games Tic-Tac-Toe and 3x3 Hex.
In Tic-Tac-Toe, which is played on a 3-by-3 grid with square cells, the objective is to get three-in-a-row horizontally, vertically, or diagonally. If the board is filled without either player having three-in-a-row, the game ends in a draw. An example game is shown in Figure 1 (top) in which the first moving player (red) wins via the diagonal.
In 3x3 Hex, which is played on a rhombus with hexagonal cells, the first moving player’s objective is to connect the top edge of the board to bottom edge, while the second moving player’s objective is to connect the left edge of the board to the right edge. Draws are not possible because the board provably cannot be filled without a player winning (gale1979game, Hex Theorem). An example game is shown in Figure 1 (bottom) in which the first moving player (red) wins by connecting the top and bottom edges.
@@
In the dark and phantom variants of these games, actions are concealed from the non-acting player, creating imperfect information. These imperfect-information variants have two rulesets—classical and abrupt—which differ in how they handle the situation in which the acting player attempts to place a piece on a cell occupied by their opponent’s piece. In the classical ruleset, the acting player must select a different cell. In abrupt, the acting player loses their turn.
@@
    OpenSpiel (openspiel19) support: DH3, PTTT, and ADH3 are already implemented in OpenSpiel, making it easy for other researchers to use them. To address APTTT’s absence, we provide our own game implementation.
@@
    Precedent: Existing work has used PTTT and ADH3 for head-to-head evaluations (mmd23; escher23; ftrl-ent23; reg-enough24), offering external evidence for their relevance.
@@
The size of these games makes it non-trivial to implement exploitability computations. To accommodate their magnitude, we utilize sequence-form representations, which are approximately 1000 times more compact than the explicit game trees, alongside dynamic payoff matrix generation to circumvent explicit matrix storage. This memory-efficient approach enables usage on commodity hardware.
@@ 5 Experiments
While our high-performance exploitability computations open the door to evaluating model-free DRL for IIGs more rigorously, the novelty of these benchmarks necessitates particular care in experimental design and interpretation. Simply showing that tuned PG methods perform “well” on these benchmarks would not suffice to evidence the policy gradient hypothesis, as there are no numbers for FP, DO, or CFR-based DRL algorithms against which to compare. The obvious recourse—which we take—is to report our own results for FP, DO, and CFR-based DRL algorithms. However, since we have hypothesized these algorithms will underperform, our role in tuning them should be carefully scrutinized. To aid readers in this endeavor, we provide detailed documentation of our algorithm implementations, hyperparameter tuning procedure, and hyperparameter tuning results, in addition to final results.
@@ 5.1 Algorithms and Implementation
We choose algorithms to evaluate using two criteria. First, that the collective set of algorithms include representatives from FP-based algorithms, DO-based algorithms, CFR-based algorithms, and generic PG methods, so as to make our experiments sufficiently comprehensive to provide evidence for or against the policy gradient hypothesis. Second, that algorithms be implemented by a reliable external source in a fashion requiring as little adaptation as possible for OpenSpiel (openspiel19) compatibility, so as to reduce the probability of an implementation error.
In accordance with these criteria, the algorithms we include are NFSP (nfsp16), PSRO (psro17), ESCHER (escher23), R-NaD (deepnash22), PPO (ppo17), PPG (ppg20), and MMD (mmd23). This selection includes one FP-based algorithm (NFSP), one DO-based algorithm (PSRO), one CFR-based algorithm (ESCHER), three generic PG methods (PPO, PPG, MMD), and one non-standard PG method (R-NaD) inspired by fforel21. Algorithm implementations were sourced from OpenSpiel (openspiel19), except PPG and ESCHER, which we sourced from CleanRL (cleanrl22) and the official ESCHER repository (escher23), respectively. These choices are summarized in Table 2.
Table 2: Algorithms and implementation sources.
Algorithm Class Implementation
NFSP FP OpenSpiel
PSRO DO OpenSpiel
ESCHER CFR Official
R-NaD Non-Standard PG OpenSpiel
PPO Generic PG OpenSpiel
PPG Generic PG CleanRL
MMD Generic PG Adapted from PPO
Of the algorithm implementation code that we wrote or modified, some of the larger potential sources of error include moving NFSP from TensorFlow (tf15) to PyTorch (pytorch19), adding multi-agent support for PPO and PPG, and implementing MMD as a modification of PPO. To corroborate the correctness of these pieces of code, we reproduce the NFSP exploitability results for Leduc Hold’em from OpenSpiel, as detailed in Section D.1, and the MMD approximate exploitability results for ADH3 and PTTT from mmd23, as detailed in Section D.2. We enable external verification of the correctness of these pieces of code, and the rest of our algorithm implementations, by open sourcing our codebase.
Figure 2: Exploitability results. For each game-algorithm pair, the box-and-whisker depicts the distribution of final exploitability over the runs from the hyperparameter tuning launch (left) and evaluation launch (right) with square-root y-axis scale. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG).
@@ 5.2 Design
For the hyperparameter tuning launch, for each combination of the 4 games and 7 algorithms, we test 50 hyperparameter configurations for 10 million steps over 3 seeds. To create these configurations, we modify the default positive real-valued hyperparameters by independent randomly sampled powers of 2—multiplying for the vast majority of hyperparameters, but exponentiating for some hyperparameters upper bounded by 1. We take the default values for hyperparameters from the implementation source in almost all cases. The most notable exceptions to this are the network architecture and optimizer, for which we impose a 3-layer 512-hidden unit fully connected network and Adam (adam15) on every algorithm, and the entropy coefficient, for which we apply the value from mmd23 to all generic PG methods. We provide a comprehensive list of hyperparameters, default values, and links to the specific lines of code of the external sources from which those default values were taken in Appendix F.
For the evaluation launch, we run the 5 strongest hyperparameter configurations from the first launch, as measured by lowest average final exploitability, with 10 fresh seeds, for 10 million steps, for each combination of the 4 games and 7 algorithms. From these evaluation runs, we report both exploitabilities and head-to-head comparisons. For the head-to-head comparisons, we select the model with the lower median final exploitability over the 10 seeds for each hyperparameter configuration, and compare these selected models for each pair of algorithms for each game.
@@ 5.3 Results
We summarize three of the main findings from our experiments: exploitability performance, head-to-head performance, and the effect of the entropy coefficient on the performance of generic PG methods.
@@
Figure 2 plots the distribution of final exploitabilities for each game and algorithm in pairs of box-and-whisker plots for the hyperparameter tuning launch (left) and the evaluation launch (right). Over these runs, ESCHER was uniformly non-competitive. PSRO was largely non-competitive, showing poor performance in ADH3, PTTT, and APTTT and having only sporadic success in DH3, for which several hyperparameter configurations found Nash equilibria on some seeds but not others. NFSP and R-NaD approached or matched the performance of generic PG methods in select cases, but typically underperformed them. The generic PG methods performed the strongest, and were roughly on par with one another.
@@
Figure 3 plots head-to-head comparisons for each pair of algorithms. Relative performance patterns are similar to those for exploitability: The generic PG methods matched or defeated other algorithms while maintaining rough parity among themselves, R-NaD either approached the performance of the generic PG methods or underperformed them, while NFSP and, especially, ESCHER and PSRO were non-competitive.
Figure 3: Head-to-head evaluations. The number in each cell is the expected utility of the row algorithm against the column algorithm, assuming each plays half of the games as the first moving player. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG), which are segregated by the dashed red line.
Entropy Coefficient Analysis
In Section 3, we suggest that one possible reason for the poor performance of generic PG methods in existing literature is that the entropy coefficients required to achieve good performance were too far removed from typical choices. To investigate, we plot exploitability as a function of entropy coefficient in Figure 4. The results are consistent with this possible explanation: Not only do unilateral changes in entropy coefficient produce drastic changes in exploitability, but the entropy coefficients with the best average performance are between 0.05 and 0.2, larger than any of the default entropy coefficients for PPO in Stable Baselines (hill2018stablebaselines), CleanRL (cleanrl22), RLlib (liang2018rllib), OpenSpiel (openspiel19), PufferLib (suarez2024pufferlibmakingreinforcementlearning), RL-Games (rl-games2021), and Tianshou (tianshou), which, as detailed in Table 4, range between 0 and 0.01.
@@ 6 Discussion
The experimental results unambiguously support the policy gradient hypothesis. Not only did the FP, DO, and CFR-based DRL algorithms fail to outperform the generic PG methods, they, alongside the non-standard PG method, were largely not even competitive. On top of that, the experiments are consistent with the possibility that poorly chosen entropy coefficients contribute to the poor performance of generic PG methods in existing literature.
@@
First, the games on which we performed experiments are homogeneous: 1) Imperfect information is introduced via hidden actions; 2) there is no chance; 3) there are only terminal non-zero rewards, which are binary; and 4) actions represent piece placements onto a 3-by-3 board. We believe that neither the relative performance of the algorithms we evaluated nor the large optimal entropy coefficients for PG methods we observed would be universally consistent across adversarial IIGs at large (indeed, relative performance is not even consistent across these homogeneous games).
@@
Third, the experimental results reflect our particular experimental conditions, including: 1) the default hyperparameter values, 2) the hyperparameters we tuned, 3) the manner in which we tuned hyperparameters, and 4) performance evaluation at 10 million steps. These conditions are indeed particular, and should not be misconstrued to represent optimal algorithm performance. It is possible that strong performing hyperparameters exist for NFSP, PSRO, ESCHER, or R-NaD, but were not found by our hyperparameter sweep because they were too many orders of magnitude away from the defaults, or because they required too many different defaults to be changed simultaneously, or because they involved different categorical hyperparameters—which we did not tune. It is also possible that the hyperparameters we used would achieve strong performance, given more training steps—a possibility consistent with preliminary experiments, which suggested that all algorithms continue to decrease exploitability beyond 10 million steps.
To power community efforts to address or investigate these second and third sets of limitations, as well as to give creators of existing and future DRL algorithms for IIGs the opportunity to demonstrate the efficacy of their algorithms, we release our exploitability computations in an accessible Python package with OpenSpiel-compatible bindings. We detail this package, called exp-a-spiel, in Appendix A. We encourage its use not only as it pertains to the policy gradient hypothesis, but also more generally for future IIGs research.
@@ 7 Conclusion
This work hypothesizes and presents evidence that generic deep PG methods are strong approaches for adversarial IIGs. Our results, along with those of mappo22, who demonstrated PPO’s effectiveness in cooperative IIGs, and mmd23, who demonstrated MMD’s effectiveness in adversarial IIGs, add to a growing body of evidence for the possibility that a single, simple PG method could serve as a universal algorithm for DRL in games. We believe this possibility is an aspiration worth pursuing.
8 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of the work, none which we feel must be specifically highlighted here.
@@ Appendix A The exp-a-spiel Package
We release a Python library to efficiently compute exploitability of policies for the four benchmark games introduced in Section 4. The library supports outputting state representations compatible with OpenSpiel (openspiel19), making it straightforward to evaluate the exploitability of policies trained using the latter library.
@@ A.1 Exploitability Computation
Internally, exp-a-spiel evaluates the exploitability of policies by carrying out computations using the so-called sequence-form representation of the four games (romanovskii; vonStengel1996; koller1996). This representation is roughly 1000 times more compact than the actual game trees, allowing us to compute exact exploitability (and head-to-head values) on commodity hardware. This representation is comprised of two objects: (i) the treeplexes $\cal X,\cal Y$ of the two players (hoda2010smoothing), and (ii) the sequence-form payoff matrix $A$ of the game. Every policy for player 1 admits an equivalent vector $x\in\cal X$ (resp. $y\in\cal Y$ for player 2), and the expected payoff corresponding to player 1 can be computed in closed form via the bilinear form $x^{\top}Ay$. Due to the size of the games, the sequence-form payoff matrix $A$ is not stored explicitly in memory by exp-a-spiel, but is rather materialized on the fly in a multi-threaded fashion as needed.

### v2

@@ 4 Benchmarks
Unfortunately, evaluating DRL algorithms in adversarial IIGs is not straightforward because of the dearth of efficient exploitability computations for large IIGs. As a result of this dearth, only a small number of works in the literature (Steinberger et al., 2020; Gruslys et al., 2020; Hu et al., 2023) make the effort to report exploitability in large games.
The rest of the literature predominantly resorts to some combination of two unsatisfactory metrics. The first is exploitability in the same small EFGs that served as benchmarks for tabular solvers, such as Leduc Hold’em (Southey et al., 2005). Benchmarking in these small games is unsatisfactory because success requires learning a precise near-Nash policy for a handful of repeatedly revisited information states—a fundamentally different challenge from large games, where success requires learning a policy that is strong (but not necessarily numerically close to Nash) for a vast number of unvisited information states.
The second is head-to-head evaluations in large games. These evaluations are unsatisfactory because the intransitive dynamics typically present in imperfect-information games, such as rock-beats-scissors-beats-paper-beats-rock, muddle interpretations of relative strength. Exacerbating the issue, there is no community coordination on head-to-head opponents, precluding direct cross-paper comparisons.
@@
To address this predicament, we set out to implement efficient computations for four large games—Phantom Tic-Tac-Toe (PTTT), 3x3 Dark Hex (DH3), Abrupt Phantom Tic-Tac-Toe (APTTT), and 3x3 Abrupt Dark Hex (ADH3)—ourselves. These games are imperfect-information twists on the perfect-information alternating-moves games Tic-Tac-Toe and 3x3 Hex.
@@
In Tic-Tac-Toe, which is played on a 3-by-3 grid with square cells, the objective is to get three-in-a-row horizontally, vertically, or diagonally. If the board is filled without either player having three-in-a-row, the game ends in a draw. An example game is shown in Figure 1 (left) in which the first moving player (red) wins via the diagonal.
In 3x3 Hex, which is played on a rhombus with hexagonal cells, the first moving player’s objective is to connect the top edge of the board to bottom edge, while the second moving player’s objective is to connect the left edge of the board to the right edge. Draws are not possible because the board provably cannot be filled without a player winning (Gale, 1979, Hex Theorem). An example game is shown in Figure 1 (right) in which the first moving player (red) wins by connecting the top and bottom edges.
In the dark and phantom variants of these games, actions are concealed from the non-acting player, creating imperfect information. These imperfect-information variants have two rulesets—classical and abrupt—which differ in how they handle the situation in which the acting player attempts to place a piece on a cell occupied by their opponent’s piece. In classical, the acting player must select a different cell. In abrupt, the acting player loses their turn.
@@
    OpenSpiel (Lanctot et al., 2019) support: DH3, PTTT, and ADH3 are already implemented in OpenSpiel, making it easy for other researchers to use them. To address APTTT’s absence, we provide our own game implementation.
@@
    Precedent: Existing work uses PTTT and ADH3 for head-to-head evaluations (Sokota et al., 2023; McAleer et al., 2023; Meng et al., 2023; Liu et al., 2024), providing external evidence of their relevance.
@@
Indeed, the size of these games makes it nontrivial to implement exploitability computations. To accommodate their magnitude, we utilize sequence-form representations, which are approximately 1000 times more compact than the explicit game trees, alongside dynamic payoff matrix generation to circumvent explicit matrix storage. This memory-efficient approach enables usage on commodity hardware.
@@ 5 Experiments
While our high-performance exploitability computations open the door to evaluating model-free DRL for IIGs more rigorously, the novelty of these benchmarks necessitates particular care in experimental design and interpretation. Simply showing that tuned PG methods perform “well” on these benchmarks would not provide strong evidence for the policy gradient hypothesis, as there are no numbers for FP-, DO-, or CFR-based DRL algorithms against which to compare. The obvious recourse—which we take—is to report our own results for FP, DO, and CFR-based DRL algorithms. However, since we have hypothesized these algorithms will underperform, our role in tuning them should be carefully scrutinized. To aid readers in this endeavor, we provide detailed documentation of our algorithm implementations, hyperparameter tuning procedure, and hyperparameter tuning results, in addition to final results.
@@ 5.1 Algorithms and Implementation
We choose algorithms to evaluate using two criteria. First, that the collective set of algorithms includes representatives from FP-based algorithms, DO-based algorithms, CFR-based algorithms, and generic PG methods, so as to make our experiments sufficiently comprehensive to provide evidence for or against the policy gradient hypothesis. Second, that algorithms be implemented by a reliable external source in a fashion requiring as little adaptation as possible for OpenSpiel (Lanctot et al., 2019) compatibility, so as to reduce the probability of an implementation error.
In accordance with these criteria, the algorithms we select are NFSP (Heinrich and Silver, 2016), PSRO (Lanctot et al., 2017), ESCHER (McAleer et al., 2023), R-NaD (Perolat et al., 2022), PPO (Schulman et al., 2017), PPG (Cobbe et al., 2021), and MMD (Sokota et al., 2023). This selection includes one FP-based algorithm (NFSP), one DO-based algorithm (PSRO), one CFR-based algorithm (ESCHER), three generic PG methods (PPO, PPG, MMD), and one non-standard PG method (R-NaD) inspired by Perolat et al. (2021). Algorithm implementations were sourced from OpenSpiel (Lanctot et al., 2019), except PPG and ESCHER, which we sourced from CleanRL (Huang et al., 2022) and the official ESCHER repository (McAleer et al., 2023), respectively.
Of the algorithm implementation code that we wrote or modified, some of the larger potential sources of error include moving NFSP from TensorFlow (Abadi et al., 2015) to PyTorch (Paszke et al., 2019), adding multi-agent support for PPO and PPG, and implementing MMD as a modification of PPO. To corroborate the correctness of these pieces of code, we reproduce the NFSP exploitability results for Leduc Hold’em from OpenSpiel, as detailed in Section D.1, and the MMD approximate exploitability results for ADH3 and PTTT from Sokota et al. (2023), as detailed in Section D.2. We enable external verification of the correctness of these pieces of code, and the rest of our algorithm implementations, by open sourcing our codebase at github.com/nathanlct/IIG-RL-Benchmark.
Figure 2: Exploitability results. For each game-algorithm pair, the box and whisker depicts the distribution of final exploitability over the runs from the hyperparameter tuning launch (left) and evaluation launch (right) with square-root y-axis scale. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG).
@@ 5.2 Design
For the hyperparameter tuning launch, for each combination of the 4 games and 7 algorithms, we test 50 hyperparameter configurations for 10 million steps over 3 seeds. To create these configurations, we modify the default positive real-valued hyperparameters by independent randomly sampled powers of 2—multiplying for the vast majority of hyperparameters, but exponentiating for some hyperparameters upper bounded by 1. We take the default values for hyperparameters from the implementation source in almost all cases. The most notable exceptions to this are the network architecture and optimizer, for which we impose a 3-layer 512-hidden unit fully connected network and Adam (Kingma and Ba, 2015) on every algorithm, and the entropy coefficient, for which we apply the value from Sokota et al. (2023) to all generic PG methods. We provide a comprehensive list of hyperparameters, default values, and links to the specific lines of code of the external sources from which those default values were taken in Appendix G.
For the evaluation launch, we run the 5 strongest hyperparameter configurations from the first launch, as measured by lowest average final exploitability, with 10 fresh seeds, for 10 million steps, for each combination of the 4 games and 7 algorithms. From these evaluation runs, we report both exploitabilities and head-to-head comparisons. For the head-to-head comparisons, we select the policy with the lower median final exploitability over the 10 seeds for each hyperparameter configuration, as well as an approximate Nash equilibrium policy computed using discounted CFR (Brown and Sandholm, 2019a), and compare these selected policies for each pair of algorithms for each game.
@@ 5.3 Results
We summarize two of the main findings from our experiments: exploitability performance and head-to-head performance.
@@
Figure 2 plots the distribution of final exploitabilities for each game and algorithm in pairs of box-and-whisker plots for the hyperparameter tuning launch (left) and the evaluation launch (right). Over these runs, ESCHER was uniformly noncompetitive. PSRO was largely noncompetitive, showing poor performance in ADH3, PTTT, and APTTT and having only sporadic success in DH3, for which several hyperparameter configurations found Nash equilibria on some seeds but not others. NFSP and R-NaD approached or matched the performance of generic PG methods in select cases, but typically underperformed them. The generic PG methods performed the strongest, and were roughly on par with one another.
@@
Figure 3 plots head-to-head comparisons for each pair of algorithms. Relative performance patterns are similar to those for exploitability: The generic PG methods matched or defeated other algorithms—and even generally outperformed the approximate Nash equilibrium against PSRO—while maintaining rough parity among themselves; R-NaD either approached the performance of the generic PG methods or underperformed them, while NFSP and, especially, ESCHER and PSRO were noncompetitive.
Figure 3: Head-to-head evaluations. The number in each cell is the expected return of the row algorithm against the column algorithm when each plays half of the games as the first moving player. R-NaD, NFSP, ESCHER, and PSRO failed to outperform generic PG methods (MMD, PPO, PPG), which are segregated by the dashed red lines.
@@ 6 Discussion
The experimental results unambiguously support the policy gradient hypothesis. Not only did the FP-, DO-, and CFR-based DRL algorithms fail to outperform the generic PG methods, they, alongside the nonstandard PG method, were largely noncompetitive.
@@
First, the games on which we performed experiments are homogeneous: 1) Imperfect information is introduced via hidden actions; 2) there is no chance; 3) there are only terminal nonzero rewards, which are binary; and 4) actions represent piece placements onto a 3-by-3 board. We do not believe that the relative performance of the algorithms we evaluated would be universally consistent across adversarial IIGs at large (indeed, it is not even consistent across these homogeneous games).
@@
Third, the experimental results reflect our particular experimental conditions, including: 1) the default hyperparameter values, 2) the hyperparameters we tuned, 3) the manner in which we tuned hyperparameters, and 4) performance evaluation at 10 million steps. These conditions are indeed particular, and should not be misconstrued to represent optimal algorithm performance. It is possible that strong performing hyperparameters exist for NFSP, PSRO, ESCHER, or R-NaD, but were not found by our hyperparameter sweep because they are too many orders of magnitude away from the defaults, or because they require too many different defaults to be changed simultaneously, or because they involve different categorical hyperparameters—which we did not tune. It is also possible that the hyperparameters we used would achieve strong performance, given more training steps—a possibility consistent with preliminary experiments, which suggested that all algorithms continue to decrease exploitability beyond 10 million steps.
To power community efforts to address or investigate these second and third sets of limitations, as well as to give creators of existing and future DRL algorithms for IIGs the opportunity to demonstrate the efficacy of their algorithms, we release our exploitability computations in a Python package with OpenSpiel-compatible bindings at github.com/gabrfarina/exp-a-spiel. We detail this package, called exp-a-spiel, in Appendix A. We encourage its use not only as it pertains to the policy gradient hypothesis, but also more generally for future IIGs research.
@@ 7 Conclusion
This work hypothesizes and presents evidence that generic deep PG methods are strong approaches for adversarial IIGs. Our results, along with those of Yu et al. (2022), who demonstrated PPO’s effectiveness in cooperative IIGs, and Sokota et al. (2023), who demonstrated MMD’s effectiveness in adversarial IIGs, add to a growing body of evidence for the possibility that a single, simple PG method could serve as a universal algorithm for DRL in games. We believe this possibility is an aspiration worth pursuing.
8 Acknowledgments
This work was supported by Award #2125858 NRT-AI: Convergent, Responsible, and Ethical Artificial Intelligence Training Experience for Roboticists; NSF #2340651; NSF #2402650; DARPA #HR00112490431; ARO W911NF-24-1-0193; the NYU IT High-Performance Computing resources, services, and staff expertise; and ONR grant #N000142212121.
We thank Jeremy Cohen and Alexander Robey for helpful feedback.
9 Erratum
arXiv version 1 claims to show results for MMD with constant reverse KL coefficients. In fact, it shows results for scheduled reverse KL coefficients that decrease over time. We thank Chang Lei for bringing this to our attention.
@@ Appendix A The exp-a-spiel Package
We release a Python library to efficiently compute exploitability of policies for the four benchmark games introduced in Section 4. The library supports outputting state representations compatible with OpenSpiel [Lanctot et al., 2019], making it straightforward to evaluate the exploitability of policies trained using the latter library.
@@ A.1 Exploitability Computation
Internally, exp-a-spiel evaluates the exploitability of policies by carrying out computations using the so-called sequence-form representation of the four games [Romanovskii, 1962, von Stengel, 1996, Koller et al., 1996]. This representation is roughly 1000 times more compact than the actual game trees, allowing us to compute exact exploitability (and head-to-head values) on commodity hardware. This representation is comprised of two objects: (i) the treeplexes $\cal X,\cal Y$ of the two players [Hoda et al., 2010], and (ii) the sequence-form payoff matrix $A$ of the game. Every policy for player 1 admits an equivalent vector $x\in\cal X$ (resp. $y\in\cal Y$ for player 2), and the expected payoff corresponding to player 1 can be computed in closed form via the bilinear form $x^{\top}Ay$. Due to the size of the games, the sequence-form payoff matrix $A$ is not stored explicitly in memory by exp-a-spiel, but is rather materialized on the fly in a multi-threaded fashion as needed.